OpenStudy (anonymous):

A triangle has side lengths of 6, 8, and 10. Is it a right triangle? Explain.

OpenStudy (anonymous):

One of the famous Pythagorean triangles, yes.

OpenStudy (anonymous):

Thank you (:

OpenStudy (solomonzelman):

\(\huge\color{ blue }{\huge {\bbox[5pt, cyan ,border:2px solid purple ]{ \bf a^2+b^2=c^2 }}}\) \(\Large\color{red}{ \bf 6^2+8^2=10^2 }\) \(\Large\color{red}{ \bf 36+64=100 }\) \(\Large\color{red}{ \bf 100=100 }\) \(\Huge\color{green }{ \bf ✓ }\)